A plano convex lens( μ=1.5) with radius of curvature 20 cm is silvered at the curved surface. The power of the system is..?


1
Expert's answer
2021-07-05T16:05:41-0400

Answer:-

Given

"\mu =1.5 \ \ , R_1=\infin \ \ \ , \\R_2=-20cm (the \ object \ lies \ on \ the \ plane \ side)"

we know that

"\frac{1}{f}=(\mu-1)(\frac{1}{R_1}-\frac{1}{R_2})"

putting the values

"\frac{1}{f}=(1.5-1)(\frac{1}{\infin}-\frac{1}{-20})"

"\frac{1}{f}=\frac{1}{40}\\
\boxed{f=40 cm}"

"D=\frac{1}{40cm}" [1 meter=100cm]


"D=\frac{100}{40}"

"\boxed{power=2.5 \ D}"